Solve by completing the square:
x2 + 8x + 13 = 0

Answer by stanbon(75887)

Solve by completing the square:
x^2 + 8x + 13 = 0
x^2+8x + ? = -13 + ?
x^2 + 8x + 4^2 = -13 + 4^2
(x+4)^2 = 3
x+4 = sqrt3 or x+4 = -sqrt3
x = -4+sqrt(3) or x = -4-sqrt(3)
===================
Cheers,
Stan H.